### v2

# M3TR: Generalist HD Map Construction with Variable Map Priors

Abstract
Autonomous vehicles require road information for their operation, usually in form of HD maps. Since offline maps eventually become outdated or may only be partially available, online HD map construction methods have been proposed to infer map information from live sensor data. A key issue remains how to exploit such partial or outdated map information as a prior. We introduce M3TR (Multi-Masking Map Transformer), a generalist approach for HD map construction both with and without map priors. We address shortcomings in ground truth generation for Argoverse 2 and nuScenes and propose the first realistic scenarios with semantically diverse map priors. Examining various query designs, we use an improved method for integrating prior map elements into a HD map construction model, increasing performance by mAP. Finally, we show that training across all prior scenarios yields a single Generalist model, whose performance is on par with previous Expert models that can handle only one specific type of map prior. M3TR thus is the first model capable of leveraging variable map priors, making it suitable for real-world deployment. https://github.com/immel-f/m3tr
1 Introduction
In order to drive safely, autonomous vehicles need to understand the geometry and topology of the roads as well as the traffic rules that apply to them. Current systems employ detailed semantic high-definition (HD) maps that provide this rich knowledge, but are primarily created using offline SLAM approaches. However, mapping the entire world in advance and maintaining HD maps to account for changes in the road layout does not scale. Therefore, recent advances in computer vision aim to perceive HD map information with onboard sensors [11, 17, 13, 14, 5, 28].
This task of online vectorized HD map construction uses sensor data, e.g. from cameras or LiDAR sensors, to detect vectorized map elements (lane markings, road borders, etc.) with their semantic meaning. Compared to offline HD planning maps however, the output of online HD map construction models still lacks a large amount of information.
Since HD maps only become outdated gradually, in addition to live sensor data, oftentimes some kind of vectorized prior map information is available and still up-to-date. This prior can range from navigation maps on a road level to parts of an HD map or even complete and recent HD maps.
Existing work that incorporates prior information falls short for three main reasons: Detection transformer queries are used to provide vectorized priors to the model [22], but the option space for query design is left underexplored. Furthermore, previous approaches focus on specialized models that can only exploit a single type of map prior that is assumed to be known in advance – an unrealistic expectation for real-world deployment. Finally, current approaches lack a task definition and evaluation metric that both uses meaningful prior scenarios and can differentiate map elements with a prior and those that need to be perceived online.
To address these points, we propose M3TR, a generalist HD map construction method with variable map priors.
Contributions
Our contributions can be summarized as follows:
We combine several recent label improvements and introduce semantically richer labels to move the HD map construction task closer towards real HD planning maps.
We systematically investigate the HD map completion task, i.e. map construction with prior map information. We introduce meaningful prior scenarios and quantify the importance of different prior element categories by evaluating their performance impact on those perceived online.
We use synthetic map masking as efficient augmentation for HD map construction with priors.
We propose a novel query design to incorporate map priors on a point query and query set level that considerably improves detection performance on the Argoverse 2 and nuScenes datasets by up to mAP.
We introduce a novel training regime which results in a single model for all map prior scenarios. This Generalist model achieves performance on par with specialized models without needing to know which kind of map information is available.
2 Related Work
Related work can be grouped into two main categories: Common online HD map construction methods without priors and methods that use prior vectorized map information.
2.1 Online HD Map Construction without Priors
Detection transformer (DETR) [4] based architectures can be used to provide vectorized map element detections, handling HD map polyline and polygon elements in their original sparse representation. To detect consistent map elements in the surrounding scene, a bird’s eye view (BEV) feature grid, representing a fixed environment area, is generated by transforming 2D image features with methods proposed in general 3D object detection and BEV segmentation [20, 12, 6]. MapTR [13, 14] implements fast detection for complete map elements by modifying the original object queries of the transformer decoder to represent polylines and polygons with a fixed number of points. This enables fast parallel transformer decoding in contrast to early autoregressive approaches like VectorMapNet [17].
Recent contributions in the online HD map construction task show two significant improvements to the MapTR and MapTRv2 baselines, concentrating on query design and formulation. The first set [28, 7] improves single-shot detection performance by utilizing complete map element shapes and masks in the detection query representation.
2.2 Online HD Map Construction with Priors
Methods discussed in the previous section take only sensor data into account. In real-world autonomous systems, maps ranging from navigation maps to HD maps are used for at least routing, extending to motion prediction, path planning, and other driving tasks. Since this map information becomes outdated piece by piece, online HD map construction that utilizes still up-to-date parts of maps as an optional prior is an attractive solution from an application perspective.
MapEX [22] was among the first to propose a detection query design allowing for both existing map element transformer queries as prior and regular learned transformer queries used for detecting unknown map elements. We use it as a baseline, but improve not only its evaluation scheme, but also the query design and model capabilities.
PriorDrive [27] proposes a HD map construction framework which integrates either SD navigation maps, incomplete HD maps or online constructed HD maps from previous drives at the same location. SMERF [19] incorporates a SD map prior by first encoding SD map elements with a transformer encoder and fusing them with the BEV feature grid via cross attention, showing improvements on the OpenLaneV2 [23] dataset detection and topology metrics. The approach of [22] to incorporate existing map prior with varying degradation levels was extended by [2] and consecutively [24] to use heavily modified map prior inputs to simulate outdated and incorrect map priors. This expands the training task to map verification, change detection, and map update, showing a significant sim-to-real gap on real public [10] or proprietary [2] data.
As mentioned in Sec. 1, previous works show a number of shortcomings, which will be discussed in the next section along with our improvements.
3 Method
In this section we describe our main contributions. Sec. 3.1 discusses the improved ground truth, Sec. 3.2 the proposed HD map completion task, Sec. 3.3 our Generalist training regime, Sec. 3.4 the map masking augmentation, and Sec. 3.5 the map prior query design.
3.1 Ground Truth Maps for Real World Autonomous Driving
Most recent HD map construction models are trained using labels that have largely been unchanged since VectorMapNet [17] despite having major shortcomings. This becomes particularly evident when compared with semantic HD planning maps used in map-based autonomous driving stacks using e.g. OpenDrive [1] or Lanelet2 [21].
First, previously used labels distinguish only three classes of map elements: boundary, divider and pedestrian crossing. They hence lack information that is necessary for autonomous driving, such as the lane topology or the distinction between dashed and solid dividers which determines if lane change maneuvers are allowed. LaneGAP [15] introduces lane centerlines to represent the topology, but does it in a way that is compatible with other map elements.
Furthermore, as pointed out and also solved by [5], the label generation algorithms of [17, 13, 14, 26, 27] produce missing or cut-off map element instances and merge actually independent elements inconsistently across frames. The third issue with current labels is the geographic overlap of the data splits pointed out by [16], leading to leakage between training and evaluation data.
To solve all described issues, we provide new ground truth labels. They complement the centerlines from [15, 14], which encode lane topology, with dashed and solid lane dividers as individual classes. For Argoverse 2, we generate labels by converting the original maps to Lanelet2 [21], which provides not only a label generation algorithm [9], but also checks for the validity of ground truth maps. Similarly, in the future, Lanelet2 could serve as common label format across datasets, forming the basis for HD map construction foundation models.
3.2 The HD Map Completion Task
Existing work that utilizes HD map priors has three open issues. Since real map changes are not labeled in ground truth maps in public datasets [10] and rare, map priors are derived synthetically. Unfortunately, previous approaches did not follow a systematic map change generation scheme with realistic assumptions [22, 27, 24]. Since real changes do not occur at random, but rather follow a local pattern with semantic correlation [2], we define so-called map prior scenarios.
| Name | Description |
| Ego lane is masked out. | |
| Ego road is masked out. | |
| Only road boundaries are provided as prior. | |
| Only lane center lines are provided as prior. | |
| No map prior. |
A map prior scenario consists of a map prior and sensor data . Map priors are derived from the complete ground truth map using the scenario specific prior generator which masks out or selects only specific map elements
All semantically meaningful scenarios are listed in Tab. 2 and partially visualized in Fig. 3. Not only do these scenarios reflect real changes more accurately, we can also investigate the importance of individual map features as a prior by evaluating their impact on detection performance for the elements that need to be perceived online.
This directly leads to the issue already pointed out by [24]: current evaluation metrics do not differentiate between map elements that are available as prior and those that need to be perceived online [22, 27, 2]. However, transformer models quickly learn to pass through prior elements almost identically and, if known as prior, any downstream application would prefer the map prior over the corresponding, but possibly noisy prediction. Hence, we propose to focus on exactly those map elements which are unknown to the model at inference time.
The standard evaluation metric for methods with vectorized output [14, 5, 17] is the mean average precision (mAP), using the Chamfer distance with thresholds of . The mAP is averaged across the average precision (AP) of the individual label classes: dashed dividers, solid dividers, road boundaries, lane centerline paths and pedestrian crossings, with the class specific AP averaged across the Chamfer distance thresholds . Analogously, we define the mean average completion precision, , which uses not the entire map , but only the map elements which are missing in the specific scenario prior .
3.3 Generalist and Expert Models
Previous works train one model for each map prior scenario which is unrealistic for deployment in real autonomous systems. All models would need to be readily available in GPU memory and the suitable model would need to be correctly selected by a not yet existing oracle that identifies the available prior category. We refer to these models as Experts and instead propose a Generalist model that can exploit arbitrary parts of HD maps as a prior. Instead of only one, the Generalist is trained on all scenarios . As we show below, while needing no extra memory or compute, the Generalist almost reaches the performance of specialized Experts. We visualize the distinction in Fig. 3.
3.4 Map Masking as Augmentation
To train a Generalist with synthetically derived map priors, various training regimes are conceivable. As depicted in Fig. 4, to derive map prior scenarios, one could naively split the dataset into equal disjoint smaller datasets and use each part to derive one kind of map prior scenario :
Instead, we propose to use the synthetic generation of prior scenario as augmentation for generic HD map construction. This means that the entire dataset is used to derive each map prior scenario and, hence, the augmented scenario set :
This exploits the entire combinatorial variety of dataset diversity and map prior categories and leads to an -fold increase in training data, promising greater generalization performance.
3.5 Query Design
In recent work, queries of the detection transformer have emerged as the main way to supply the model with prior information [22, 24, 27, 2, 18, 5]. How exactly these queries are composed and where they are inserted is often neglected and not described in detail. BEV detection transformer queries consist of different sub-elements and map prior queries can be composed in many different ways and inserted at multiple points, making the available option space quite large. We explore that option space to incorporate prior map knowledge on two architectural levels, the point query design and the query set design, and use MapEX [22] as our baseline approach.
Point Query Design
For each map element that is to be predicted, a fixed set of points is used as map decoder queries. Each point query consists of two vectors which are concatenated: the point embedding and the positional embedding .
[…]
Query Set Design
MapTRv2 [14] proposed one-to-many (O2M) matching, a source of significant performance gains compared to the original MapTR one-to-one (O2O) matching. We explore two possible ways to adapt it to map prior information which are depicted in Fig. 5.
In the (Single Map Prior) query set design only a single repetition of queries makes use of map prior queries while auxiliary queries are purely learned, just like in the original MapTRv2. In contrast, the (Multiple Map Prior) query set design includes map prior information in a tiling fashion, once for every repetition of the ground truth. This allows the incorporation of map prior knowledge in the auxiliary queries as well.
We follow MapEX [22] for the loss, including the pre-attribution of map prior instances during the Hungarian assignment, which we extend to the tiled map prior queries. Outside of instances related to map priors, the MapEX loss is equivalent to the loss of the MapTRv2 base architecture. A more detailed description of the pre-attribution can be found in the supplementary material.
4 Experiments
We conduct experiments on the Argoverse 2 and nuScenes datasets to validate our method, with Argoverse 2 as the main dataset. Sec. 4.1 elaborates on the choices of dataset and metric, Sec. 4.2 on the implementation and Sec. 4.3 discusses the performance of M3TR in comparison with existing baselines.
4.1 Dataset and Metric
As mentioned above, we validate our method on the Argoverse 2 [25] and nuScenes datasets [3], the standard public datasets for HD map construction. Both Argoverse 2 and nuScenes contain 1000 driving sequences, covering 17 km² and 5 km², respectively [16]. Since nuScenes has only 40,000 samples compared to 158,000 for Argoverse 2, contrary to most existing work, we regard Argoverse 2 as our primary dataset for evaluation. As discussed in Sec. 3.1, we use a novel kind of ground truth that resolves a number of problems compared to the labels used in [13, 14, 17, 26].
As our metric, we use the mean average completion precision defined in Sec. 3.2 to compare the investigated methods.
To simulate a real use case with various priors, groups of expert models are compared with a single generalist model, calculating the mean for each class across prior scenarios. This assumes for the benefit of the experts that a perfect oracle for prior-to-model assignment exists and that no mixing of prior categories occurs.
4.2 Implementation Details and Baseline
[…]
All models are trained until convergence, i.e. on 24 / 110 epochs for experts and 54 / 224 epochs for generalist models on Argoverse 2 / nuScenes respectively, and the best checkpoint is shown. The generalist models were trained on nine map prior scenarios for Argoverse 2 and seven for nuScenes. The four scenarios not explicitly shown are missing only centerlines / pedestrian crossings / road borders / dividers.
4.3 Map Completion Performance
| Dataset: Argoverse 2 | = AP for Masked Elements only | Var. Prior w/o Oracle | ||||||||
| Method | Map Prior | vs. [22] | ||||||||
| MapTRv2 [14] | 37.9 | 55.0 | 49.7 | 48.2 | 41.7 | 46.5 | †+0.0 | - | - | |
| MapEX* [22] Models | 45.3 | 64.5 | 53.4 | 52.8 | 44.9 | 52.2 | - | ✗ | ||
| 41.5 | 62.4 | 54.9 | 55.3 | 45.5 | 51.9 | - | ||||
| 37.7 | 56.0 | - | 50.6 | 44.5 | 47.2 | - | ||||
| 43.2 | 61.8 | 58.1 | - | 42.8 | 51.5 | - | ||||
| †37.9 | †55.0 | †49.7 | †48.2 | †41.7 | †46.5 | - | ||||
[…]
| Mean | 41.1 | 59.9 | 54.0 | 51.7 | 43.9 | 49.9 | - | |||
| M3TR Expert Models | 51.7 | 69.4 | 56.3 | 55.4 | 49.7 | 56.5 | +4.3 | ✗ | ||
| 44.8 | 66.5 | 57.0 | 57.8 | 48.7 | 55.0 | +3.1 | ||||
| 40.2 | 57.3 | - | 54.7 | 49.2 | 50.2 | +3.0 | ||||
| 45.1 | 63.2 | 61.1 | - | 48.6 | 55.0 | +3.5 | ||||
| †37.9 | †55.0 | †49.7 | †48.2 | †41.7 | †46.5 | †+0.0 | ||||
[…]
| Mean | 43.9 | 62.3 | 56.0 | 54.0 | 47.5 | 52.6 | +2.7 | |||
| M3TR Generalist | 48.8 | 67.8 | 59.5 | 54.8 | 51.8 | 56.5 | +4.3 | ✓ | ||
| 45.7 | 64.4 | 57.0 | 56.9 | 51.1 | 55.0 | +3.1 | ||||
| 41.2 | 57.3 | - | 53.0 | 48.0 | 49.9 | +2.7 | ||||
| 42.5 | 59.3 | 57.4 | - | 45.6 | 51.2 | -0.3 | ||||
| 40.4 | 55.4 | 50.3 | 49.4 | 43.9 | 47.9 | +1.4 | ||||
[…]
| Mean | 43.7 | 60.8 | 56.0 | 53.5 | 48.1 | 52.1 | +2.2 | |||
| vs. [22] | |||||||
| 49.4 | 69.3 | 56.9 | 54.8 | 49.7 | 56.0 | +3.8 | |
| 41.4 | 65.6 | 55.8 | 56.2 | 48.6 | 54.4 | +2.5 | |
| 42.7 | 58.4 | - | 52.7 | 46.4 | 49.7 | +2.7 | |
| 42.7 | 59.9 | 55.4 | - | 43.6 | 50.4 | -1.1 | |
| 40.5 | 56.0 | 49.0 | 48.4 | 41.8 | 47.2 | +0.7 | |
| Mean | 43.3 | 61.8 | 54.3 | 53.0 | 46.0 | 51.5 | +1.6 |
| Map Query Enc. | |||||||
| Point Enc. | O2MMMP | dsh. | sol. | bou. | cen. | ped. | |
| A [22] | — | 52.2 | 45.3 | 64.5 | 53.4 | 52.8 | 44.9 |
| B | — | 52.4 (+0.2) | 46.8 | 65.3 | 53.0 | 52.5 | 44.5 |
| C | — | 53.5 (+1.3) | 48.4 | 66.7 | 55.4 | 50.6 | 46.5 |
| C | 56.5 (+4.3) | 51.7 | 69.4 | 56.3 | 55.4 | 49.7 | |
As we view Argoverse 2 as our main dataset for evaluation, we investigate more map prior scenarios on it than on nuScenes. We first discuss the results on Argoverse 2 along with ablations on the map query encoder and the map masking as augmentation. Then we present the slightly reduced set of experiments on the nuScenes dataset.
Results on Argoverse 2
Tab. 3 shows the performance of the M3TR expert and generalist variants as well as a MapEX expert baseline for five selected map prior scenarios on the Argoverse 2 dataset. All methods using map priors show enhanced average precision compared to the prior-less scenario, with varying benefit depending on the supplied map prior. A qualitative example of the varying benefit of different map priors can be seen in Fig. 6.
For almost all scenarios, the M3TR experts variants substantially improve the prediction performance compared to the MapEX baseline. Except for the scenario, the Generalist model matches the performance of the M3TR Expert models in their expert scenarios as well.
In the more real-world use case with varying map priors, the generalist likewise outperforms the baseline models in the average of all scenarios, only using a fifth of the VRAM and without an oracle for perfect prior scenario to expert assignment. Such a scenario to expert assignment system does not exist yet and provides another substantial obstacle for real-world use of groups of expert models.
The generalist model additionally shows better performance in the no map prior scenario (), equivalent to the standard pure online HD map construction task, without architectural changes relevant for this scenario compared to the MapTRv2 base. We conjecture that the various prior scenarios function as augmentation that also helps the model learn HD map construction without any prior.
Ablations on Argoverse 2
[…]
Results on nuScenes
| Dataset: nuScenes | = AP for Masked Elements only | Var. Prior w/o Oracle | ||||||||
| Method | Map Prior | vs. [22] | ||||||||
| MapTRv2 [14] | 12.5 | 19.1 | 32.4 | 29.1 | 21.6 | 22.9 | †+0.0 | - | - | |
| MapEX* [22] Models | 13.2 | 21.1 | - | 31.0 | 22.0 | 21.9 | - | ✗ | ||
| 16.6 | 26.0 | 39.8 | - | 23.4 | 26.4 | - | ||||
| †12.5 | †19.1 | †32.4 | †29.1 | †21.6 | †22.9 | - | ||||
[…]
| Mean | 14.1 | 22.1 | 36.1 | 30.1 | 22.3 | 23.7 | - | |||
| M3TR Expert Models | 15.3 | 26.7 | - | 34.9 | 28.3 | 26.3 | +4.4 | ✗ | ||
| 23.1 | 33.2 | 46.6 | - | 27.8 | 32.5 | +6.1 | ||||
| †12.5 | †19.1 | †32.4 | †29.1 | †21.6 | †22.9 | †+0.0 | ||||
[…]
| Mean | 17.0 | 26.3 | 39.5 | 32.0 | 25.9 | 27.2 | +3.5 | |||
| M3TR Generalist | 14.5 | 23.2 | - | 32.7 | 24.7 | 23.8 | +1.9 | ✓ | ||
| 15.3 | 24.4 | 38.6 | - | 24.4 | 25.7 | -0.7 | ||||
| 12.4 | 20.0 | 31.8 | 29.8 | 23.4 | 23.5 | +0.6 | ||||
[…]
| Mean | 14.1 | 22.5 | 35.2 | 31.3 | 24.2 | 24.3 | +0.5 | |||
Table 6 shows results on the nuScenes dataset on a reduced set of map prior scenarios. Compared to the baseline, the increase in expert performance is even larger than on Argoverse 2. While the generalist model improves upon the baseline it shows reduced performance compared to M3TR experts.
Notably, the M3TR generalist model still shows increased performance for the case without prior compared to the MapTRv2 base, underlining once more the effectiveness of map masking as augmentation.
[…]
5 Conclusion
This work proposes M3TR, a novel generalist approach for HD map construction with variable map priors.
We introduce improved ground truth that moves the HD map construction task closer towards real HD planning maps. Using it, we define the HD map completion task, including a systematic set of map prior scenarios combined with a metric that focuses on those elements not given as a map prior. Building on our proposed ground truth as a common label format, we plan to extend our approach towards foundational map models that can generalize not only across tasks, but also across datasets.
[…]
Training with partially masked out maps not only allows utilizing prior map information, but also serves as augmentation for HD map construction without any map prior.
Finally, our novel Generalist training regime yields a single model that can exploit variable map priors and still reaches the performance of specialized Experts. At the same time, it requires only a constant amount of memory and does not need to know which kind of map information is available. This makes M3TR the first model for HD map construction with priors that can be deployed in the real world.
Supplementary Material
6 Remaining Details of Novel Ground Truth
This section presents more details of our novel ground truth, namely a qualitative example of reprojected commonly used and new labels and a comparison of label features on the nuScenes dataset.
Qualitative Comparison
Fig. 8 contrasts commonly used ground truth and our new ground truth for the same sample in Argoverse 2, reprojected into the camera images. Fig. 8(a) with the commonly used ground truth illustrates that many lane dividers are not included in the labels, such as the left dashed divider of the vehicles ego lane and the solid divider in front of the vehicle separating the two driving directions. Additionally, the lack of distinction between solid and dashed dividers leaves crucial planning information for the vehicle missing. Fig. 8(b), our proposed new ground truth, contains these features and, as seen in the reprojection, provides a much more comprehensive collection of road information.
Labels on nuScenes
The nuScenes dataset has a different and less detailed map format compared to Argoverse 2 and as a result not all label features from Tab. 1 are applicable. Tab. 7 provides a similar overview for the nuScenes dataset. nuScenes does not include 3D map labels, making this aspect of comparison moot. Furthermore, the original label generation code of VectorMapNet [17] was designed for nuScenes, therefore the divider artifacts from Argoverse 2 are not present. It is important to note though that the original nuScenes maps themselves have missing divider annotations, for example intersection areas do not have annotated lane dividers at all [3]. Like for Argoverse 2, M3TR also separates lane dividers by type and uses lane centerlines for nuScenes. The code to generate lane centerline labels was taken from MapTRv2 [14], as it already exists in the codebase, however no model trained with these labels is mentioned or evaluated in the MapTRv2 paper.
7 Details of Loss Pre-Attribution
As mentioned in Sec. 3.5, we follow MapEX [22] for the training loss, including the pre-attribution of map prior instances during assignment. A visualization of the pre-attribution strategy can be seen in Fig. 7. All ground truth instances given to the model as a prior are also included in the loss term, so that the model is trained to also predict ground truth elements, receiving a complete map as output. As a result, the model learns very quickly to pass through prior map elements for our task of map completion.
To help the model in identifying queries with map priors, MapEX proposes to remove instances of these queries from the regular Hungarian assignment and directly assign them to the respective ground truth elements, as it is already known for the initial map prior queries which ground truth instance they belong to.
8 Additional Evaluation Results and Details
| Dataset: Argoverse 2 | AP = AP for All Elements | |||||||
| Method | Map Prior | mAP | vs. [22] | |||||
| MapTRv2 [14] | 37.9 | 55.0 | 49.7 | 48.2 | 41.7 | 46.5 | - | |
| MapEX* [22] Models | 77.4 | 85.5 | 97.1 | 79.3 | 58.9 | 79.7 | - | |
| 67.3 | 74.6 | 94.7 | 70.5 | 54.4 | 72.3 | - | ||
| †37.7 | †56.0 | 97.1 | †50.6 | †44.5 | 57.2 | - | ||
| †43.2 | †61.8 | †58.1 | 95.7 | †42.8 | 60.3 | - | ||
[…]
| Mean | 56.4 | 69.5 | 86.8 | 74.0 | 50.2 | 67.4 | - | |
| M3TR Expert Models | 80.5 | 87.6 | 97.1 | 80.6 | 61.0 | 81.4 | +1.7 | |
| 69.4 | 77.2 | 94.7 | 72.7 | 55.4 | 73.9 | +1.6 | ||
| †40.2 | †57.3 | 99.9 | †54.7 | †49.2 | 60.3 | +3.1 | ||
| †45.1 | †63.2 | †61.1 | 96.3 | †48.6 | 62.9 | +2.6 | ||
[…]
| Mean | 58.8 | 71.3 | 88.2 | 76.1 | 53.6 | 69.6 | +2.3 | |
| M3TR Generalist | 78.9 | 86.8 | 97.6 | 80.3 | 63.3 | 81.4 | +1.7 | |
| 69.5 | 75.8 | 95.0 | 72.6 | 57.7 | 74.1 | +1.8 | ||
| †41.2 | †57.3 | 99.9 | †53.0 | †48.0 | 59.9 | +2.7 | ||
| †42.5 | †59.3 | †57.4 | 96.9 | †45.6 | 60.3 | +0.0 | ||
[…]
| Mean | 58.0 | 69.8 | 87.5 | 75.7 | 53.7 | 68.9 | +1.6 | |
This section presents additional evaluation results on Argoverse 2 and qualitative examples to complement Sec. 4.
Remaining Implementation Details
As shown in Fig. 5, the total number of O2M queries is chosen as a multiple of the ground truth repetitions to match with the tiling. In total all models have O2O and O2M queries, therefore having times the amount of O2O queries as O2M queries. MapTRv2 [14] uses a fixed number of 300 O2M queries in their original implementation.
Argoverse 2 AP with Prior
For Argoverse 2, we also report the average precision for all elements in the ground truth in Tab. 8, whether they were included in the prior or not.
In conjunction with the from Tab. 3, this can give an overview on the total accuracy of the resulting map from the model depending on the supplied prior. For elements in the prior, the AP is calculated with the output of the model for those elements, not using the original ground truth prior. Additionally, we can also evaluate the ability to pass through given prior elements. In an application setting, one would in any case include known map elements in the constructed HD map from the prior directly to have these elements be guaranteed correct.
All models reproduce the prior almost perfectly, with M3TR showing slightly better metrics for the and priors. The M3TR Experts and Generalist also still outperform the baseline in regular AP, mirroring Sec. 4, though less pronounced for the and scenarios. This can be attributed to the fact that the prior, which both methods pass through very well, is now included in the metric. The share of online perceived instances, where the performance differences lie, is therefore diluted in the regular AP evaluated here.
Qualitative Results
In the following, we show qualitative results of M3TR together with MapTRv2 [14] and MapEX [22] on the Argoverse 2 [25] and nuScenes [3] datasets.
Fig. 9 displays the M3TR Generalist and MapTRv2 on Argoverse 2, both with no prior. With our masking as augmentation training regime M3TR shows increased performance in complex scenes, without any architectural changes.
Fig. 10 and Fig. 11 demonstrate the M3TR Generalist and the respective MapEX Experts on every evaluated prior scenario and show the prediction of M3TR with no prior as well. The M3TR Generalist is able to use the prior to better perceive the missing online elements, with higher accuracy compared to the MapEX baseline and only using a single model. As seen in Fig. 10, we further observed that MapEX sometimes inserts predictions that are logically inconsistent with the given prior, e.g. road borders that cross a lane centerline, hinting at a lessened understanding of the prior compared to M3TR.
For nuScenes, Fig. 12 presents a comparison of the M3TR and MapEX Experts, with the no prior expert also shown for reference. The general performance is decreased of all models compared to Argoverse 2 with large errors without any prior. In conjunction, the performance gain when supplying a prior is stronger. The mentioned tendency of MapEX to sometimes make predictions that are logically inconsistent with the given prior can also be seen here.

### v3

@@
-M3TR: Generalist HD Map Construction with Variable Map Priors
+M3TR: A Generalist Model for Real-World HD Map Completion
 Abstract
-Autonomous vehicles require road information for their operation, usually in form of HD maps.
-Since offline maps eventually become outdated or may only be partially available, online HD map construction methods have been proposed to infer map information from live sensor data.
-A key issue remains how to exploit such partial or outdated map information as a prior.
-We introduce M3TR (Multi-Masking Map Transformer), a generalist approach for HD map construction both with and without map priors.
-We address shortcomings in ground truth generation for Argoverse 2 and nuScenes and propose the first realistic scenarios with semantically diverse map priors.
-Examining various query designs, we use an improved method for integrating prior map elements into a HD map construction model, increasing performance by mAP.
-Finally, we show that training across all prior scenarios yields a single Generalist model, whose performance is on par with previous Expert models that can handle only one specific type of map prior.
-M3TR thus is the first model capable of leveraging variable map priors, making it suitable for real-world deployment.
+Autonomous vehicles rely on HD maps for their operation, but offline HD maps eventually become outdated.
+For this reason, online HD map construction methods use live sensor data to infer map information instead.
+Research on real map changes shows that oftentimes entire parts of an HD map remain unchanged and can be used as a prior.
+We therefore introduce M3TR (Multi-Masking Map Transformer), a generalist approach for HD map completion both with and without offline HD map priors.
+As a necessary foundation, we address shortcomings in ground truth labels for Argoverse 2 and nuScenes and propose the first comprehensive benchmark for HD map completion.
+Unlike existing models that specialize in a single kind of map change, which is unrealistic for deployment, our Generalist model handles all kinds of changes, matching the effectiveness of Expert models.
+With our map masking as augmentation regime, we can even achieve a mAP improvement without a prior.
+Finally, by fully utilizing prior HD map elements and optimizing query designs, M3TR outperforms existing methods by mAP while being the first real-world deployable model for offline HD map priors.
 https://github.com/immel-f/m3tr
@@
 Current systems employ detailed semantic high-definition (HD) maps that provide this rich knowledge, but are primarily created using offline SLAM approaches.
-However, mapping the entire world in advance and maintaining HD maps to account for changes in the road layout does not scale.
+However, maintaining such offline HD maps to account for changes in no time is infeasible.
 Therefore, recent advances in computer vision aim to perceive HD map information with onboard sensors [11, 17, 13, 14, 5, 29].
@@
 Compared to offline HD planning maps however, the output of online HD map construction models still lacks a large amount of information.
-Since HD maps only become outdated gradually, in addition to live sensor data, oftentimes some kind of vectorized prior map information is available and still up-to-date.
-This prior can range from navigation maps on a road level to parts of an HD map or even complete and recent HD maps.
-Existing work that incorporates prior information falls short for three main reasons: Detection transformer queries are used to provide vectorized priors to the model [22], but the option space for query design is left underexplored.
-Furthermore, previous approaches focus on specialized models that can only exploit a single type of map prior that is assumed to be known in advance – an unrealistic expectation for real-world deployment.
-Finally, current approaches lack a task definition and evaluation metric that both uses meaningful prior scenarios and can differentiate map elements with a prior and those that need to be perceived online.
-To address these points, we propose M3TR, a generalist HD map construction method with variable map priors.
+Recent research [10] showed that maps only gradually become outdated and that oftentimes some parts of vectorized (offline) map information is still up-to-date and could be used as prior.
+Concretely, only specific, often semantically coherent elements are invalidated, while leaving the rest unchanged.
+This leads to the situation where a map perception model needs to fill in the invalid parts using the remaining HD map and online sensor information, a task which we refer to as HD map completion.
+Existing work that incorporates prior information falls short for three main reasons: While detection transformer queries are used to provide vectorized priors to the model [23], they fail to fully utilize all prior map information.
+Furthermore, current approaches lack a clear task definition and evaluation metric that can differentiate prior map elements and those that need to be perceived online.
+Finally and most importantly, previous models specialize on a single kind of map prior that is assumed to be known in advance.
+Since any part of an offline HD map could change, this is an unrealistic expectation for real-world deployment.
 Contributions
-Our contributions can be summarized as follows:
-We combine several recent label improvements and introduce semantically richer labels to move the HD map construction task closer towards real HD planning maps.
-We systematically investigate the HD map completion task, i.e.
-map construction with prior map information.
-We introduce meaningful prior scenarios and quantify the importance of different prior element categories by evaluating their performance impact on those perceived online.
-We use synthetic map masking as efficient augmentation for HD map construction with priors.
+To address these points, we present M3TR (Multi-Masking Map Transformer), a generalist HD map completion model with the following contributions:
+A new HD map completion benchmark for models with prior offline HD map information.
+This includes semantically richer labels and the first metric that explicitly focuses on the performance for elements without a prior.
 We propose a novel query design to incorporate map priors on a point query and query set level that considerably improves detection performance on the Argoverse 2 and nuScenes datasets by up to mAP.
-We introduce a novel training regime which results in a single model for all map prior scenarios.
-This Generalist model achieves performance on par with specialized models without needing to know which kind of map information is available.
+We introduce a novel training regime which yields a single model that can make use of any HD map prior.
+This Generalist model achieves performance on par with specialized models without needing to know which kind of map information is available, even improving performance without a prior by up to 1.4 mAP.
 2 Related Work
@@
 In real-world autonomous systems, maps ranging from navigation maps to HD maps are used for at least routing, extending to motion prediction, path planning, and other driving tasks.
-Since this map information becomes outdated piece by piece, online HD map construction that utilizes still up-to-date parts of maps as an optional prior is an attractive solution from an application perspective.
+Since this map becomes outdated piece by piece, online HD map construction that utilizes still up-to-date parts of maps as an optional prior is an attractive solution from an application perspective.
 MapEX [23] was among the first to propose a detection query design allowing for both existing map element transformer queries as prior and regular learned transformer queries used for detecting unknown map elements.
@@
 As mentioned in Sec. 1, previous works show a number of shortcomings, which will be discussed in the next section along with our improvements.
-3 Method
-In this section we describe our main contributions.
-Sec.
-3.1 discusses the improved ground truth, Sec.
-3.2 the proposed HD map completion task, Sec.
-3.3 our Generalist training regime, Sec.
-3.4 the map masking augmentation, and Sec.
-3.5 the map prior query design.
-3.1 Ground Truth Maps for Real World Autonomous Driving
+3 The HD Map Completion Benchmark
+In this section we describe the novel HD map completion benchmark. Sec. 3.1 discusses our improved ground truth while Sec. 3.2 presents the proposed HD map completion task.
+3.1 Improved Ground Truth Maps for Real World Autonomous Driving
 Most recent HD map construction models are trained using labels that have largely been unchanged since VectorMapNet [17] despite having major shortcomings.
-This becomes particularly evident when compared with semantic HD planning maps used in map-based autonomous driving stacks using e.g.
-OpenDrive [1] or Lanelet2 [21].
-First, previously used labels distinguish only three classes of map elements: boundary, divider and pedestrian crossing.
-They hence lack information that is necessary for autonomous driving, such as the lane topology or the distinction between dashed and solid dividers which determines if lane change maneuvers are allowed.
-LaneGAP [15] introduces lane centerlines to represent the topology, but does it in a way that is compatible with other map elements.
-Furthermore, as pointed out and also solved by [5], the label generation algorithms of [17, 13, 14, 26, 27] produce missing or cut-off map element instances and merge actually independent elements inconsistently across frames.
-The third issue with current labels is the geographic overlap of the data splits pointed out by [16], leading to leakage between training and evaluation data.
-To solve all described issues, we provide new ground truth labels.
-They complement the centerlines from [15, 14], which encode lane topology, with dashed and solid lane dividers as individual classes.
-For Argoverse 2, we generate labels by converting the original maps to Lanelet2 [21], which provides not only a label generation algorithm [9], but also checks for the validity of ground truth maps.
-Similarly, in the future, Lanelet2 could serve as common label format across datasets, forming the basis for HD map construction foundation models.
+The labels lack information that is necessary for autonomous driving, issues in the label generation algorithms introduce errors into the ground truth instances and a geographic overlap leads to leakage between training and evaluation data.
+Fixes to these issues have been proposed in different works [14, 15, 5, 16], however they are scattered and not united in one single ground truth set.
+Hence, as foundation for the proposed HD map completion benchmark, we combine these improvements into one label set together with a novel separation of dashed and solid dividers.
+A comparison with previously used labels is listed in Tab.
+1.
+For a detailed description with qualitative examples we refer to the supplementary material.
 3.2 The HD Map Completion Task
-Existing work that utilizes HD map priors has three open issues.
-Since real map changes are not labeled in ground truth maps in public datasets [10] and rare, map priors are derived synthetically.
-Unfortunately, previous approaches did not follow a systematic map change generation scheme with realistic assumptions [22, 27, 24].
-Since real changes do not occur at random, but rather follow a local pattern with semantic correlation [2], we define so-called map prior scenarios.
+Our focus in this work is on using offline HD maps as priors that became outdated and thus partially invalid.
+This idea is based on the largest public dataset of real map changes, Trust but Verify [10].
+When applying existing work to outdated offline HD maps, three open issues arise: Map changes in public datasets are not labeled on a point level and rare, requiring map priors to be derived synthetically [10].
+Unfortunately, the map change generation schemes of previous approaches [23, 28, 25] follow assumptions that are not applicable for outdated offline HD maps.
+Real changes do not occur at random, but rather follow a local pattern with semantic correlation [2], that only affects specific elements and leaves most elements unchanged.
+This was also observed in [10], where changes remove, modify or add semantically coherent elements rather than randomly drop/add elements or apply noise.
 | Name | Description |
@@
 | No map prior. |
-A map prior scenario consists of a map prior and sensor data .
-Map priors are derived from the complete ground truth map using the scenario specific prior generator which masks out or selects only specific map elements
-All semantically meaningful scenarios are listed in Tab.
-2 and partially visualized in Fig.
-3.
-Not only do these scenarios reflect real changes more accurately, we can also investigate the importance of individual map features as a prior by evaluating their impact on detection performance for the elements that need to be perceived online.
-This directly leads to the issue already pointed out by [24]: current evaluation metrics do not differentiate between map elements that are available as prior and those that need to be perceived online [22, 27, 2].
-However, transformer models quickly learn to pass through prior elements almost identically and, if known as prior, any downstream application would prefer the map prior over the corresponding, but possibly noisy prediction.
+To better align our synthetic priors with real priors, we define adapted map prior scenarios consisting of a map prior and sensor data .
+Map priors are derived from the complete ground truth map using the scenario specific prior generator which masks out or selects only specific map elements:
+The task of the model is to reconstruct the complete map from the given partial prior and the sensor information.
+MapEX [23] has already begun moving in this direction by including as a scenario, however the other scenarios include modifications like point-level noise that are not applicable for outdated offline HD maps, but only for maps perceived in previous time steps.
+We show in Sec.
+5 that the semantic class of map prior has a strong influence on the model performance and therefore propose to separate map prior scenarios semantically.
+This enables a systematic investigation to guide future efforts in data collection or map maintenance.
+The prior scenarios are listed in Tab.
+2 and visualized in Fig.
+5.
+Fig.
+2 shows that real map changes [10] can easily be categorized into the proposed scenarios.
+Fig.
+2(a) has the own lane become blocked, resulting in invalidated elements akin to .
+In Fig.
+2(b), a bike lane is added that causes the ego road to become invalid, similar to .
+The scenarios assume that it is known beforehand which elements are no longer valid, a task for which separate proposed solutions exist [25, 20].
+In turn however, this also brings a large benefit: These map prior scenarios avoid a cause of the significant sim-to-real gap already noted for artificial map changes [2], namely the fact that most synthetic changes are not logically consistent with the sensor data.
+This is because the reconstruction task is indifferent to whether elements are masked synthetically or if elements become masked due to real map changes, as both mask semantically coherent elements.
+3.3 A Prior-Aware HD Map Completion Metric
+To measure map completion performance, we need to solve an issue already pointed out by [25]: current evaluation metrics do not differentiate between map elements that are available as prior and those that need to be perceived online [23, 28, 2].
+However, transformer models quickly learn to pass through prior elements almost identically and, if known as prior, any downstream application would prefer the map prior over the corresponding, possibly noisy prediction.
 Hence, we propose to focus on exactly those map elements which are unknown to the model at inference time.
@@
 The mAP is averaged across the average precision (AP) of the individual label classes: dashed dividers, solid dividers, road boundaries, lane centerline paths and pedestrian crossings, with the class specific AP averaged across the Chamfer distance thresholds .
-Analogously, we define the mean average completion precision, , which uses not the entire map , but only the map elements which are missing in the specific scenario prior .
-3.3 Generalist and Expert Models
-Previous works train one model for each map prior scenario which is unrealistic for deployment in real autonomous systems.
-All models would need to be readily available in GPU memory and the suitable model would need to be correctly selected by a not yet existing oracle that identifies the available prior category.
-We refer to these models as Experts and instead propose a Generalist model that can exploit arbitrary parts of HD maps as a prior.
-Instead of only one, the Generalist is trained on all scenarios .
-As we show below, while needing no extra memory or compute, the Generalist almost reaches the performance of specialized Experts.
-We visualize the distinction in Fig.
-3.
-3.4 Map Masking as Augmentation
-To train a Generalist with synthetically derived map priors, various training regimes are conceivable.
-As depicted in Fig.
-4, to derive map prior scenarios, one could naively split the dataset into equal disjoint smaller datasets and use each part to derive one kind of map prior scenario :
-Instead, we propose to use the synthetic generation of prior scenario as augmentation for generic HD map construction.
-This means that the entire dataset is used to derive each map prior scenario and, hence, the augmented scenario set :
-This exploits the entire combinatorial variety of dataset diversity and map prior categories and leads to an -fold increase in training data, promising greater generalization performance.
+Analogously, to evaluate completion performance, we define the mean average completion precision, , which uses not the entire map , but only the map elements which are missing in the specific scenario.
+4 A Deployable HD Map Completion Model
+This section describes the novel M3TR (Multi-Masking Map Transformer) model itself.
+Sec.
+4.2 presents our Generalist training regime, Sec.
+4.3 how to use map masking as augmentation, and Sec.
+4.1 the novel map prior query design.
 4.1 Query Design
 In recent work, queries of the detection transformer have emerged as the main way to supply the model with prior information [23, 25, 28, 2, 18, 5].
-How exactly these queries are composed and where they are inserted is often neglected and not described in detail.
 BEV detection transformer queries consist of different sub-elements and map prior queries can be composed in many different ways and inserted at multiple points, making the available option space quite large.
@@
 We explore two possible ways to adapt it to map prior information which are depicted in Fig.
-5.
+3.
 In the (Single Map Prior) query set design only a single repetition of queries makes use of map prior queries while auxiliary queries are purely learned, just like in the original MapTRv2.
@@
 A more detailed description of the pre-attribution can be found in the supplementary material.
+4.2 Generalist and Expert Models
+Previous works train one model for each map prior scenario which is unrealistic for deployment in real autonomous systems.
+All models would need to be readily available in GPU memory and the suitable model would need to be correctly selected by a not yet existing oracle that identifies the available prior category.
+We refer to these models as Experts and instead propose a Generalist model that can exploit arbitrary parts of HD maps as a prior.
+Instead of only one, the Generalist is trained on all scenarios .
+As we show below, while needing no extra memory or compute, it is on par with specialized Experts.
+We visualize the distinction in Fig.
+5.
+4.3 Map Masking as Augmentation
+To train a Generalist with synthetically derived map priors, various training regimes are conceivable.
+As depicted in Fig.
+4, to derive map prior scenarios, one could naively split the dataset into equal disjoint smaller datasets and use each part to derive one kind of prior scenario :
+Instead, we propose to use the synthetic generation of prior scenario as augmentation for generic HD map construction.
+This means that the entire dataset is used to derive each map prior scenario and, hence, the augmented scenario set :
+This exploits the entire combinatorial variety of dataset diversity and map prior categories and leads to an -fold increase in training data, promising greater generalization performance.
 5 Experiments
@@
 5.1 Dataset and Metric
-As mentioned above, we validate our method on the Argoverse 2 [25] and nuScenes datasets [3], the standard public datasets for HD map construction.
 Both Argoverse 2 and nuScenes contain 1000 driving sequences, covering 17 km² and 5 km², respectively [16].
@@
 3.1, we use a novel kind of ground truth that resolves a number of problems compared to the labels used in [13, 14, 17, 27].
-As our metric, we use the mean average completion precision defined in Sec. 3.2 to compare the investigated methods.
-To simulate a real use case with various priors, groups of expert models are compared with a single generalist model, calculating the mean for each class across prior scenarios.
-This assumes for the benefit of the experts that a perfect oracle for prior-to-model assignment exists and that no mixing of prior categories occurs.
-4.2 Implementation Details and Baseline
-[…]
-All models are trained until convergence, i.e.
-on 24 / 110 epochs for experts and 54 / 224 epochs for generalist models on Argoverse 2 / nuScenes respectively, and the best checkpoint is shown.
-The generalist models were trained on nine map prior scenarios for Argoverse 2 and seven for nuScenes.
-The four scenarios not explicitly shown are missing only centerlines / pedestrian crossings / road borders / dividers.
-4.3 Map Completion Performance
+As our metric, we use the mean average completion precision defined in Sec. 3.2 to compare the methods.
 | Dataset: Argoverse 2 | = AP for Masked Elements only | Var. Prior w/o Oracle | ||||||||
@@
 | C | 56.5 (+4.3) | 51.7 | 69.4 | 56.3 | 55.4 | 49.7 | |
-As we view Argoverse 2 as our main dataset for evaluation, we investigate more map prior scenarios on it than on nuScenes.
-We first discuss the results on Argoverse 2 along with ablations on the map query encoder and the map masking as augmentation.
-Then we present the slightly reduced set of experiments on the nuScenes dataset.
-Results on Argoverse 2
-Tab.
-3 shows the performance of the M3TR expert and generalist variants as well as a MapEX expert baseline for five selected map prior scenarios on the Argoverse 2 dataset.
-All methods using map priors show enhanced average precision compared to the prior-less scenario, with varying benefit depending on the supplied map prior.
-A qualitative example of the varying benefit of different map priors can be seen in Fig.
-6.
-For almost all scenarios, the M3TR experts variants substantially improve the prediction performance compared to the MapEX baseline.
-Except for the scenario, the Generalist model matches the performance of the M3TR Expert models in their expert scenarios as well.
-In the more real-world use case with varying map priors, the generalist likewise outperforms the baseline models in the average of all scenarios, only using a fifth of the VRAM and without an oracle for perfect prior scenario to expert assignment.
-Such a scenario to expert assignment system does not exist yet and provides another substantial obstacle for real-world use of groups of expert models.
-The generalist model additionally shows better performance in the no map prior scenario (), equivalent to the standard pure online HD map construction task, without architectural changes relevant for this scenario compared to the MapTRv2 base.
-We conjecture that the various prior scenarios function as augmentation that also helps the model learn HD map construction without any prior.
-Ablations on Argoverse 2
-[…]
-[…]
-Results on nuScenes
+To simulate a real use case with various priors, groups of expert models are compared with a single generalist model, calculating the mean for each class across prior scenarios.
+This assumes for the benefit of the experts that a perfect oracle for prior-to-model assignment exists and that no mixing of prior categories occurs.
 | Dataset: nuScenes | = AP for Masked Elements only | Var. Prior w/o Oracle | ||||||||
@@
 | Mean | 14.1 | 22.5 | 35.2 | 31.3 | 24.2 | 24.3 | +0.5 | |||
+5.2 Implementation Details and Baseline
+[…]
+All models are trained until convergence, i.e.
+for 24 / 110 epochs for experts and 54 / 224 epochs for the generalist on Argoverse 2 / nuScenes respectively, with the best checkpoint shown.
+The generalist was trained on nine map prior scenarios for Argoverse 2 and seven for nuScenes.
+The four scenarios not explicitly shown are missing only centerlines / pedestrian crossings / road borders / dividers.
+5.3 Map Completion Performance
+As we view Argoverse 2 as our main dataset for evaluation, we investigate more map prior scenarios on it than on nuScenes.
+We first discuss the results on Argoverse 2 along with ablations on the map query encoder and the map masking as augmentation.
+Then we present the slightly reduced set of experiments on the nuScenes dataset.
+Results on Argoverse 2
+Tab.
+3 shows the performance of the M3TR expert and generalist variants as well as a MapEX expert baseline for five selected map prior scenarios on the Argoverse 2 dataset.
+All methods using map priors show enhanced average precision compared to the prior-less scenario, with varying benefit depending on the supplied map prior.
+A qualitative example of this varying benefit can be seen in Fig.
+6.
+For almost all scenarios, the M3TR experts variants substantially improve the prediction performance compared to the MapEX baseline.
+Except for the scenario, the generalist model matches the performance of the M3TR Expert models in their expert scenarios as well.
+In the more real-world use case with varying map priors, the generalist likewise outperforms the baseline models in the average of all scenarios, only using a fifth of the VRAM and without an oracle for perfect prior scenario to expert assignment.
+Such a scenario to expert assignment system does not exist yet and provides another substantial obstacle for real-world use of groups of expert models.
+The generalist model additionally shows better performance in the no map prior scenario (), equivalent to the standard pure online HD map construction task, without architectural changes relevant for this scenario compared to the MapTRv2 base.
+We conjecture that the various prior scenarios function as augmentation that also helps the model learn HD map construction without any prior.
+Ablations on Argoverse 2
+[…]
+[…]
+Results on nuScenes
 Table 6 shows results on the nuScenes dataset on a reduced set of map prior scenarios.
 Compared to the baseline, the increase in expert performance is even larger than on Argoverse 2.
-While the generalist model improves upon the baseline it shows reduced performance compared to M3TR experts.
+While the Generalist improves upon the baseline, it shows reduced performance compared to M3TR experts.
 Notably, the M3TR generalist model still shows increased performance for the case without prior compared to the MapTRv2 base, underlining once more the effectiveness of map masking as augmentation.
@@
 We introduce improved ground truth that moves the HD map construction task closer towards real HD planning maps.
-Using it, we define the HD map completion task, including a systematic set of map prior scenarios combined with a metric that focuses on those elements not given as a map prior.
-Building on our proposed ground truth as a common label format, we plan to extend our approach towards foundational map models that can generalize not only across tasks, but also across datasets.
+Using it, we define a new HD map completion benchmark, including a systematic set of prior scenarios for outdated HD maps, combined with a metric that focuses on the elements not given as a map prior.
 […]
-Training with partially masked out maps not only allows utilizing prior map information, but also serves as augmentation for HD map construction without any map prior.
-Finally, our novel Generalist training regime yields a single model that can exploit variable map priors and still reaches the performance of specialized Experts.
-At the same time, it requires only a constant amount of memory and does not need to know which kind of map information is available.
-This makes M3TR the first model for HD map construction with priors that can be deployed in the real world.
+We show that training with partially masked out maps not only allows using prior map information, but also serves as augmentation for HD map construction without any prior.
+Finally, our novel Generalist is a single model that can handle all map prior scenarios while matching the performance of specialized Experts.
+Contrary to an ensemble of experts, it requires only constant memory and does not need to know which type of map information is available.
+This makes M3TR the first real-world deployable model for HD map construction with offline HD map priors.
 Supplementary Material
 7 Remaining Details of Novel Ground Truth
-This section presents more details of our novel ground truth, namely a qualitative example of reprojected commonly used and new labels and a comparison of label features on the nuScenes dataset.
+This section presents more details of our novel ground truth, namely a more detailed description of the changes and their motivation, qualitative examples of commonly used and new labels and a comparison of label features on the nuScenes dataset.
+Detailed Description of Changes and Motivation
+Compared with semantic HD planning maps used in map-based autonomous driving stacks using e.g.
+OpenDrive [1] or Lanelet2 [22], the shortcomings of commonly used labels become particularly evident.
+We describe these shortcomings, already mentioned in Sec.
+3.1, our changes and their motivation in more detail here.
+Previously used labels distinguish only three classes of map elements: boundary, divider and pedestrian crossing.
+They hence lack information that is necessary for autonomous driving, such as the lane topology or the distinction between dashed and solid dividers which determines if lane change maneuvers are allowed.
+Recognizing this issue, LaneGAP [15] introduces lane centerlines to represent the topology, but does it in a way that is compatible with other map elements.
+Furthermore, as pointed out and also solved by [5], the label generation algorithms of [17, 13, 14, 27, 28] produce missing or cut-off map element instances and merge actually independent elements inconsistently across frames.
+The third issue with current labels is the geographic overlap of the data splits pointed out by [16], leading to leakage between training and evaluation data.
+To solve all described issues, we provide new ground truth labels.
+We complement the centerlines from [15, 14], which encode lane topology, with dashed and solid lane dividers as individual classes.
+For Argoverse 2, we generate labels by converting the original maps to Lanelet2 [22], which provides not only a label generation algorithm [9], but also checks for the validity of ground truth maps.
+Similarly, in the future, Lanelet2 could serve as common label format across datasets, forming the basis for HD map construction foundation models.
 Qualitative Comparison
+Fig.
+7 shows examples of commonly used labels and our ground truth.
+The new ground truth is semantically richer, distinguishes dashed and solid lane dividers and fixes the many missing or incorrect divider instances.
 Fig.
@@
 Fig.
-8(a) with the commonly used ground truth illustrates that many lane dividers are not included in the labels, such as the left dashed divider of the vehicles ego lane and the solid divider in front of the vehicle separating the two driving directions.
+9(a) with the commonly used ground truth illustrates that many lane dividers are not included in the labels, such as the left dashed divider of the vehicles ego lane and the solid divider in front of the vehicle separating the two driving directions.
 Additionally, the lack of distinction between solid and dashed dividers leaves crucial planning information for the vehicle missing.
 Fig.
-8(b), our proposed new ground truth, contains these features and, as seen in the reprojection, provides a much more comprehensive collection of road information.
+9(b), our proposed new ground truth, contains these features and, as seen in the reprojection, provides a much more comprehensive collection of road information.
 Labels on nuScenes
@@
 As mentioned in Sec.
-3.5, we follow MapEX [22] for the training loss, including the pre-attribution of map prior instances during assignment.
+4.1, we follow MapEX [23] for the training loss, including the pre-attribution of map prior instances during assignment.
 A visualization of the pre-attribution strategy can be seen in Fig.
-7.
+8.
 All ground truth instances given to the model as a prior are also included in the loss term, so that the model is trained to also predict ground truth elements, receiving a complete map as output.
@@
 As shown in Fig.
-5, the total number of O2M queries is chosen as a multiple of the ground truth repetitions to match with the tiling.
+3, the total number of O2M queries is chosen as a multiple of the ground truth repetitions to match with the tiling.
 In total all models have O2O and O2M queries, therefore having times the amount of O2O queries as O2M queries.
@@
 The M3TR Experts and Generalist also still outperform the baseline in regular AP, mirroring Sec.
-4, though less pronounced for the and scenarios.
+5, though less pronounced for the and scenarios.
 This can be attributed to the fact that the prior, which both methods pass through very well, is now included in the metric.
@@
 As seen in Fig.
-10, we further observed that MapEX sometimes inserts predictions that are logically inconsistent with the given prior, e.g.
+11, we further observed that MapEX sometimes inserts predictions that are logically inconsistent with the given prior, e.g.
 road borders that cross a lane centerline, hinting at a lessened understanding of the prior compared to M3TR.